OpenStudy (anonymous):

Is the equation true, false or open? 4y+8=6y+3

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Because 4y+8==>y=2 6y+3==>y=1/2 2 doesn't equal 1/2 hehe